OpenStudy (sainath):

What is an example that shows how Experimental Probability and Theoretical Probability are different? Real life example.

OpenStudy (sainath):

Would this work? Someone would think that out of 50 times of rolling a normal dice, they would get "6" at least once. But really, they got 0/50.

OpenStudy (sainath):

@jim_thompson5910

OpenStudy (sainath):

@Lord_Box @zepdrix @Zarkon

jimthompson5910 (jim_thompson5910):

Let's say you flip a coin 100 times and you record these results number of heads: 58 number of tails: 42 the experimental probability of getting heads is 58/100 = 0.58 the theoretical probability of getting heads is 1/2 = 0.50

OpenStudy (sainath):

oooh i get it

jimthompson5910 (jim_thompson5910):

the theoretical probability is the ideal target to aim for. If you flipped the coin 1000 times, you'll probably get roughly half to be heads and half to be tails. It won't be 100% perfect but it will be close. The experimental probability will get closer and closer to the theoretical probability as the sample size (n) gets larger

OpenStudy (unklerhaukus):

If you roll a single die once. You can theoretically expect to get the average value i.e. (1+2+3+4+5+6) / 6 = 21/6 = 3.5 Experimentally, you can not roll a 3.5 you can only roll a 1, 2, 3, 4, 5 or 6

OpenStudy (sainath):

oh ok. thanks dude i get it now

OpenStudy (unklerhaukus):

Another example: You go to the casino, All games at the casino are slightly in favour of the house, so theoretically, you have to expect that you will loose money. However, experimentally (if you are lucky) you can win.

OpenStudy (sainath):

ok :D ty dude